OpenStudy (anonymous):

20 pounds of a salt water solution contains 15 percent salt ,how much water must be added to weaken thesolution to 12percent salt?

OpenStudy (shadowfiend):

So let's figure out first how much salt is in the water. If we have 20 pounds of salt water, and 15 percent of it is salt, we can find the amount of salt in the water by saying: $$ 0.15 * 20 = 3 $$ So 3 pounds are salt and 17 pounds are water. Then we need 12 percent salt. We can set up an equality: $$\frac{3}{w} = \frac{12}{100}$$

OpenStudy (shadowfiend):

There, \(w\) is the amount of salt water (i.e., the total). We're saying 12 out of 100 is the same as 3 pounds out of how many total pounds. 12 out of 100 is because 12% is defined as 12 parts out of every 100. Does that make sense? Is that enough for you to find the answer?

OpenStudy (anonymous):

thank you it does :)

OpenStudy (shadowfiend):

Awesome! :)

OpenStudy (anonymous):

:)